OpenStudy (anonymous):

what is the answer to-2y – 5 = -9

OpenStudy (anonymous):

y=2

OpenStudy (anonymous):

y=2

OpenStudy (anonymous):

-2y-5=-9 -2y=-4 -y=-2 y=2